OpenStudy (anonymous):

giving medal for answer Write a formula for the sum of the measures of the angles of an n-gon (n sides). a. s = 360(n + 2) c. s = 180(n - 2) b. s = 360(n - 2) d. s = 180(n + 2)

OpenStudy (anonymous):

sum of interior angles formula for polygon is = 180(n-2) where n is the number of vértices . For example Heptagon n=7 so: 180(7-2)=180*5=900 to find the size of each angle just devide this by 7: 900/7=128.6°

OpenStudy (anonymous):

It's: s = 180(n - 2) because you can divide a polygon of "n" sides (where n > 3) into "n - 2" triangles, and the sum of the angles in any triangle is 180.

OpenStudy (anonymous):

For example with a pentagon, you will have 3 triangle (5-2)

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @leahhbe